SYMPLECTIC STRUCTURES AND DIFFEOMORPHISMS

BOGUSŁAW HAJDUK

Abstract

I discuss the problem of existence of symplectic structures on tori with different differential structures and related question of calculating the action of the group of diffeomorphisms on the homotopy classes of almost complex structures. I show that diffeomorphisms which are topologically isotopic to the identity may act nontrivially.

Mathematical Institute University of Wrocław plac Grunwaldzki 2/4 50–384 Wrocław POLAND e-mail: hajduk@math.uni.wroc.pl